OpenStudy (anonymous):

HC)Rhett is showing his work in simplifying (−7.9 + 5.2) − 1.4. Step 1: (5.2 − 7.9) − 1.4 Step 2: 5.2 − (7.9 + 1.4) Step 3: 5.2 − (9.3) Step 4: −4.1 Which statement is true? In step 1, Rhett used the distributive property. In step 2, Rhett used the associative property. In step 3, Rhett used the commutative property. In step 4, Rhett's answer is incorrect. Plz assist

OpenStudy (anonymous):

anyone

OpenStudy (anonymous):

Distributive property is if you have the following: 5(x+2) = 5x+10 Commutative property is if you have something like a x b = b x a, you really just rearrange the variables/ integers. Associative property would be when they want you to group things/ rearrange them, (a+b) +c = a+(b+c)